OpenStudy (anonymous):

A right circular cylinder has a radius of 8 and a height of pie squared. If a cube has the same volume as the cylinder, what is the length of an edge of the cube?

OpenStudy (anonymous):

first, what is the volume of the cylinder?

OpenStudy (anonymous):

the volume of the cylinder is \(\large V=\pi\cdot r^2\cdot h \), where r is the radius and h is the height...

OpenStudy (anonymous):

yes. and the formula of a c is v=a^3

OpenStudy (anonymous):

the radius is 8 and h is pie squared. So the formula is :

OpenStudy (anonymous):

ok... so what's the volume of the cylinder with the given dimensions?

OpenStudy (anonymous):

\[V=\pi(8)^{2}(\pi)^{2}\]

OpenStudy (anonymous):

So it becomes \[V=64(\pi)^{3}\] right?

OpenStudy (anonymous):

P.S. answer choices are given (A) \[4\sqrt{\pi}\] (B)\[8\sqrt{\pi}\] (C) \[4\pi \sqrt{\pi}\] (D) 4pie (E) 8pie

OpenStudy (anonymous):

correct... so now with your formula for the volume of a cube, \(\large s^3=64\pi^3 \) you'll need to solve for s....

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so how do I exactly do that?

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

got the answer. It's 4pi, D